pen 15 22 11896 | Use this mini-case to answer the following question(s). The Delux Nut Company produces a deluxe mix of almonds, cashews, peanuts and walnuts. The deluxe mix must contain at least 10% of each kind of nut. At least half of the mix must consist of almonds and cashews and it can contain at most 20% peanuts. The company also produces a companion mix of cashews, walnuts and contain at least 10% cashews, at least 30% walnuts, and between 20 and 40 are available in unlimited supply at a purchasing cost of R3,60 per kilogram. The quantities of nuts available and the purchasing costs in rand per kilogram are as follows: (Total marks 100) raisins. This mix must % raisins. Raisins Nut Almonds Cashews Peanuts Walnuts Purchasing costs in rand 12,00 10,60 2.80 8.40 Kilogram available 400 200 600 300 Both mixes are sold in 100 g packages. The company must produce at least 4 000 packages of the deluxe mix and at least 5 000 packages of the companion mix. (a) Formulate an LP model for this problem (b) Solve the LP using LINGO or SOLVER (c) Write 20 marks) (20 marks down the optimal purchasing plan and the associated purchasing

media%2Fd4c%2Fd4cdfa92-9833-4f9d-9040-c3
Answer #1

A, B, C

The boxes in green, are decision variables
The boxes in blue are constrains
The boxes in orange is the objective function

| Book41 - Excel Sign in File Home Insert Page Layout Formulas Review View Help Tell me what you want to do Connections Flash Fill - Remove DuplicatesRelationships Group Data Analysis 目Ungroup , -ー ?, Solver Show Queries Clear From Table Get Extern New Data Query Sort Fiterppy Refresh What-f Forecast Edit Links Subtotal Query Lo Recent Sources Al Columns Data Validation Analysis-sheet Get & Transform Connections Sort & Filter Data Tools Solver Parameters C21 fSUMPRODUCT(C17:G17,C20:G20) Set objective SCS21 Decision variable Production in packs Almonds Cashews Peanuts Walnuts Ralsins in one pack Total Deluxe 1.00 1.00 1.00 1.00 3 Mix By Changing Variable Cells: SBS3:SFS3,5B$11,SDS11,SFS11:SGS11 25.00% | 25.00% | 25.00% | 25.00% 10% 50 00% 10% Almond +Cashews Subject to the Constraints: BS11integer SBS11 SBS13 SBS3 integer SBS3-SBSS SCS17:SFS17 SCS19:$FS19 20% Add 11 n mix 33.33% 33.33% 33.33% SDS12-SOS14 SFS12.SGS12SF514:SGS14 SGS14-SGS16 HS11 SIS11 Beset All ad/Save Total onsilabiliy irai Make Unconstrained Variables Non-Negative Select a Solving GRG N onlinear Cost per 10.6 8.4 3.6 20 KG Total cost 56.4 Objective function Minimize Select the GRG Nonlinear engine for Solver Problems that are smooth nonlinear. Select the LP Simplex engine for linear Sover Problems, and select the Evolutionary engine for Solver problems that are non-smooth. 23 24 25 26 Sheet1

The formulation:

Decision variable g in one pack
Production in packs Almonds Cashews Peanuts Walnuts Raisins Total
Deluxe Mix 4000 12.4999975552772 37.5 20.0000012223595 30.0000012223633 =SUM(C3:G3) = 100
>= =C3/$H$3 =D3/$H$3 =E3/$H$3 =F3/$H$3
4000 >= >= >= >=
0.1 0.1 0.1 0.1
Almond + Cashews =SUM(C3:D3)/H3 <=
>= 0.2
0.5
Companion mix 5000 10 29.9999995653828 60.0000004346173 =SUM(C11:G11) = 100
>= =D11/$H$11 =F11/$H$11 =G11/$H$11
5000 >= >= >=
0.1 0.3 0.2
<=
0.4
Total =$B$3*C3+$B$11*C11 =$B$3*D3+$B$11*D11 =$B$3*E3+$B$11*E11 =$B$3*F3+$B$11*F11 =$B$3*G3+$B$11*G11
Availability constraint (g) <= <= <= <=
400000 200000 600000 300000
Cost per KG 12 10.6 2.8 8.4 3.6
Total cost =SUMPRODUCT(C17:G17,C20:G20)
Objective function Minimize

The solution:

Decision variable g in one pack
Production in packs Almonds Cashews Peanuts Walnuts Raisins Total
Deluxe Mix 4000 12.50 37.50 20.00 30.00 100 = 100
>= 12.50% 37.50% 20.00% 30.00%
4000 >= >= >= >=
10% 10% 10% 10%
Almond + Cashews 50.00% <=
>= 20%
50%
Companion mix 5000 10 30 60 100 = 100
>= 10.00% 30.00% 60.00%
5000 >= >= >=
10% 30% 20%
<=
40%
Total 50000 200000 80000 270000 300000
Availability constraint (g) <= <= <= <=
400000 200000 600000 300000
Cost per KG 12 10.6 2.8 8.4 3.6
Total cost 6292000
Objective function Minimize

D.

It does not change the solution, as the required for Almond is 50 Kg and for peanuts is 80 Kg.

So, the cost remains the same and so does the mix.

E.

It is of no use, as only 270 Kg is used, hence there is already 30 kg unused, hence 50 kg extra is of no use. Optimal solution persists.

F.

So we change the constraint from 4000 to 4500.

So the solution changes: As below

Decision variable g in one pack
Production in packs Almonds Cashews Peanuts Walnuts Raisins Total
Deluxe Mix 4500 16.67 33.33 20.00 30.00 100 = 100
>= 16.67% 33.33% 20.00% 30.00%
4500 >= >= >= >=
10% 10% 10% 10%
Almond + Cashews 50.00% <=
>= 20%
50%
Companion mix 5000 10 30 60 100 = 100
>= 10.00% 30.00% 60.00%
5000 >= >= >=
10% 30% 20%
<=
40%
Total 75000.0 200000 90000.01 285000 300000
Availability constraint (g) <= <= <= <=
400000 200000 600000 300000
Cost per KG 12 10.6 2.8 8.4 3.6
Total cost 6746000
Objective function Minimize